40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jun 3 22:48:59 2018, johannes, Update, PSL, aux laser replacement 
    Reply  Mon Jun 4 02:52:52 2018, johannes, Update, PSL, aux laser replacement 
    Reply  Mon Jun 4 11:00:37 2018, gautam, Update, PSL, aux laser replacement 
       Reply  Tue Jun 5 02:06:59 2018, gautam, Update, PSL, aux laser first (NULL) results Close-But-no.jpg
Message ID: 13911     Entry time: Sun Jun 3 22:48:59 2018     Reply to this: 13912   13913
Author: johannes 
Type: Update 
Category: PSL 
Subject: aux laser replacement 

I brought the NPRO from the Crackle experiment over to the 40m Lab and set it up on the PSL table to replace the slowly dying AUX laser. I also brought along a Faraday isolator, broadband EOM, and an ISOMET AOM with driver electronics from the optics storage in the Crackle Lab.

This laser is a much newer model, made in 2008, and still has all its mojo, but we should probably keep up the practice of turning it off when it's not going to be used for a while. I measured 320 mW leaving the laser, and 299mW of that going through the Faraday isolator, whose Brewster-angle polarizer I had to clean because they were a little dusty. While the laser output is going strong, the controller displays a power output of only 10 mW, which makes me think that the power monitoring PD is busted. This is a completely different failure mode from what we've seen with the other NPROs that we can hopefully get repaired at some point, particularly because the laser is newer, but for now it's installed on the PSL table. This likely means that the noise eater isn't working on this unit either, for different reasons, but at least we have plenty of optical power.

The setup is very similar to before, with the addition of a Faraday isolator and a broadband EOM, in case we decide to get more bandwidth in the PLL. I changed the Crystal Technologies 3200-113 200 MHz AOM for an ISOMET 80 MHz AOM with RF driver from the Crackle lab's optics storage and sized the AUX beam to a diameter of 200 micron. I couldn't locate an appropriate heat sink for the driver, which is still in factory condiction, but since the PSL AOM also runs on 80MHz I used that one instead. The two AOMs saturate at different RF powers, so care must be taken to not drive the AUX AOM too high. At 600 mV input to the driver the deflection into the first order was maximal at 73 % of the input power, with the second order beam and the first order on the other side cleary visible.

In order to speed things up I didn't spend too much time on mode-matching, but the advantage of the fiber setup is that we can always improve later if need be without affecting things downstream. I coupled the first order beam into the fiber to the AS table with 58% efficiency, and restored the beat with the PSL laser on the NewFocus 1611. The contrast there is only about 20%, netting a -20 dBm beat note. This is only a marginal improvement from before, so the PLL will work as usual, but if we get the visibility up a little in the future we won't need to amplify the PD signal for the PLL anymore.

Some more things I wanted to do but didn't get to today are

  • Measure intensity noise of aux laser
  • Measure rise and fall times of new AOM
  • Get PLL back up and running
  • Place 90/10 beamsplitter in AS path and couple IFO output into fiber (= couple fiber output into IFO)

I'll resume this work tomorrow. I turned the aux laser and the AOM driver input off. For the PSL beat the AOM drive is not needed, and the power in the optical fiber should not exceed 100 mW, so the offset voltage to the AOM RF driver has to remain below 300 mV.

ELOG V3.1.3-